OpenStudy (anonymous):

Given the system of equations presented here: 3x + 5y = 29 x + 4y = 16 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated?

OpenStudy (anonymous):

Multiply the second equation by −1 to get −x − 4y = −16 Multiply the second equation by −3 to get −3x − 12y = −48 Multiply the first equation by −1 to get −3x − 5y = −29 Multiply the first equation by −3 to get −9x − 15y = −87

OpenStudy (anonymous):

@imqwerty i got c

imqwerty (imqwerty):

no how did u get c?

OpenStudy (anonymous):

oh i see i misread it

OpenStudy (anonymous):

b ?

imqwerty (imqwerty):

yes correct :)